OpenStudy (anonymous):

HELP AWARDING METAL!! Find all solutions to the equation. sin2x + sin x = 0

Nnesha (nnesha):

do-you-know-sin-cos-properties-reciprocal-functions?? sin^2=??

OpenStudy (anonymous):

tbh, no.

OpenStudy (henriquerb):

sin (2x) = sinxcosx + cosxsinx = 2sinxcosx

OpenStudy (anonymous):

thanks :)

Nnesha (nnesha):

i-guess-2-is-exponent-right?

OpenStudy (henriquerb):

sin(x)(1+2cosx). So, sin(x)(1+2cosx)=0 when sin(x)=0 or 1+2cosx=0

OpenStudy (anonymous):

That's it?

OpenStudy (anonymous):

yep

OpenStudy (henriquerb):

Well, do you think you can handle from there?

OpenStudy (anonymous):

No not really?

OpenStudy (henriquerb):

Ok, so 2cosx = -1 and sinx=0

OpenStudy (anonymous):

yea?

OpenStudy (henriquerb):

Now you have to find the x on those equations.

OpenStudy (henriquerb):

sin x = 0 is easy, just npi, with n integer.

OpenStudy (anonymous):

Oh, ok

OpenStudy (henriquerb):

2cosx=-1 so cosx=-1/2 so x = 2pi/3 + 2pi n, with n integer or x=4pi/3 + 2pi n, with n integer

OpenStudy (anonymous):

Oh, so that's it?

OpenStudy (henriquerb):

I think that's it. If i didn't make any mistakes.

OpenStudy (anonymous):

Wow, your like a genius.

OpenStudy (anonymous):

Thank you :::)

OpenStudy (henriquerb):

I'm surely not. You're welcome. =)

OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 2 sin2x = sin x